1. A solution is prepared that contains 5g of NaCl in 200mLs of H2O. What percentage is this?

2. The molecular or gram equivalent weight for NaOH is 40. How many grams of NaOH is needed to make a 0.25 molar solution?

3. A technologist must add 20 microliters of an expensive reagent diluent to each of 25 tubes for a serial dilution. How much diluent will she need with out having extra that must be thrown out?
Answer #1

1.) Mass of NaCl = 5g

Volume of water = 200ml

density of water = 1g/ml

mass of water = 200ml/1g/ml = 200g

mass of solution = mass of NaCl + mass of water = 205g

Percentage of NaCl(w/w) = 5*100/205 = 2.44%

Percentage (w/V) = 5 *100/200 = 2.5%

2.) Molar mass of NaOH = 40g/mol

Molarity = number of moles of solute / volume of Solution (in liter)

Let volume be one liter

0.25 = n /1

moles = 0.25

Number of moles = Mass / Molar Mass
mass = moles ^ Molar Mass
Mass = 0.25 * 40 = 10 gram

3.) If one Test tube Reqires 20 microliters and there are 25 test tubes

Then 25 test tubes need 25 *20 microliters = 5* 10^-4 L = 0.5ml